The length and the breadth of a rectengle are increased by 12% and x% respectively,thereby the area of the rectengle increased by 40%. What is the value of x?
Answer by stanbon(75887)

The length and the breadth of a rectengle are increased by 12% and x% respectively,thereby the area of the rectengle increased by 40%. What is the value of x?
-----
Original DATA: length = L ; Breadth = B ; Area = LB?
----
New Area Equation:
1.12L * (1+x/100)B = 0.40LB
----
1.12(1+x/100) = 0.4
1+(x/100) = 1.25
x/100 = 0.25
x = 25
---------
Cheers,
Stan H.